OpenStudy (anonymous):

How is it wrong? \[\large 1 = \sqrt{1} = \sqrt{i^4} = i^2 = -1\]

Directrix (directrix):

The product property of square roots applies only to positive, real numbers.

OpenStudy (anonymous):

I guess it's because \(i^4 = (i^2)^2\), so we have \(\sqrt{(-1)^2}\neq\sqrt{-1}\sqrt{-1}\). Am I correct?

OpenStudy (acxbox22):

no directrix is always right....

OpenStudy (anonymous):

So I am wrong? @acxbox22

OpenStudy (anonymous):

We basically said the same thing lol...

OpenStudy (acxbox22):

well @Directrix stated a rue of exponents  because you can only use exponent rules with real numbers i is imaginary and hence is not real